Circular No. 6184

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1629–49 D. J. McKay, M. H. Wieringa, and N. P. F. McKay, The Australia

Telescope (AT) National Facility, CSIRO, report the results of AT Compact Array observations of the area of the γ-ray outburst reported on IAUC 6182: “Observations on July 6.2, 6.6, and 7.2 UT at frequencies 1380, 2378, 4800, and 8640 MHz indicate that there is no new bright radio source in the field. We have remeasured the source PMN J1628–4936 (nominated in IAUC 6182 as a possible radio counterpart): flux-density = 1.2 ± 0.1 Jy at all but 8640 MHz, where it is strongly resolved; α = 16^{h}28^{m}57^{s}.8 ± 0^{s}.2, δ = −49^{o}36′26′′.4 ± 1′′.0 (equinox 2000.0); angular size 15′′. This source is also identified with a source catalogued in 1979 as an H II region (G334.714) and with IRAS 16251–4929. On the basis of this information, we believe that the PMN object is an H II region not identified with the γ-ray object. We detect no radio source at the optical position reported by McNaught (IAUC 6183) to a limit of 10 mJy at 2378 MHz on July 6.2 and 6.6.”

R. H. McNaught now retracts his note on IAUC 6183, saying that his optical candidate for GRO J1629–49 is present on the SERC I survey (but not the ESO R). G. Garradd further reports that inspection of a UKST film taken on July 7 UT by K. S. Russell shows nothing new that is obvious down to mag ∼ 15, using the I survey photograph as a comparison; he looked a further 0^{o}.5 outward from McNaught’s original search area.

EXO 053109–6609.2 K. Dennerl, F. Haberl, and W. Pietsch, Max-Planck-Institut für Ex-

traterrestrische Physik, Garching, communicate: “We report the discovery of coherent 13.7-s pulsations from EXO 053109–6609.2, a Be transient in the Large Magellanic Cloud, with the ROSAT Position Sensitive Proportional Counter (PSPC). The pulsations are present in several observations covering intrinsic source luminosities from 1 × 10^{35} to 3 × 10^{36} erg s^{−1} (E = 0.1–2.4 keV). We determined a barycentric pulse period of 13.67133 ± 0.00005 s at 1991 Oct. 31.158 UT and a rate of period change of (+1.5 ± 0.1) × 10^{−8} during 3.3 days, which exceeds the long-term trend by at least one order of magnitude. Under the assumption that the observed period change is predominantly caused by orbital Doppler shifts, the orbital period can be constrained to the range from 4 to 40 days, with 25.4 days as the most probable value.”

1995 July 7 Daniel W. E. Green